# Revisiting Locally Differentially Private Protocols: Towards Better Trade-offs in Privacy, Utility, and Attack Resistance

### v2

@@ Abstract
Local Differential Privacy (LDP) offers strong privacy protection, especially in settings in which the server collecting the data is untrusted. However, designing LDP mechanisms that achieve an optimal trade-off between privacy, utility and robustness to adversarial inference attacks remains challenging. In this work, we introduce a general multi-objective optimization framework for refining LDP protocols, enabling the joint optimization of privacy and utility under various adversarial settings. While our framework is flexible to accommodate multiple privacy and security attacks as well as utility metrics, in this paper, we specifically optimize for Attacker Success Rate (ASR) under data reconstruction attack as a concrete measure of privacy leakage and Mean Squared Error (MSE) as a measure of utility. More precisely, we systematically revisit these trade-offs by analyzing eight state-of-the-art LDP protocols and proposing refined counterparts that leverage tailored optimization techniques. Experimental results demonstrate that our proposed adaptive mechanisms consistently outperform their non-adaptive counterparts, achieving substantial reductions in ASR while preserving utility, and pushing closer to the ASR-MSE Pareto frontier. By bridging the gap between theoretical guarantees and real-world vulnerabilities, our framework enables modular and context-aware deployment of LDP mechanisms with tunable privacy-utility trade-offs.
@@ I Introduction
The local DP model has gained widespread adoption, with major technology companies integrating it into their systems to enhance user privacy. Notable examples include Google Chrome [3] and Windows 10 [4], in which LDP protocols have been used to collect statistics while safeguarding individual data. A fundamental task under LDP guarantees is frequency estimation, which forms the basis of many advanced data analysis tasks like heavy hitter estimation [5, 6, 7], frequency monitoring [3, 4, 8, 9], multidimensional queries [10, 11, 12, 13], frequent item-set mining [14, 15] and spatial density estimation [16, 17].
@@
Due to its importance, numerous LDP frequency estimation protocols have been proposed [18], namely, Generalized Randomized Response (GRR) [19], Subset Selection (SS) [20, 21], Symmetric Unary Encoding (SUE) [3], Optimized Unary Encoding (OUE) [22], Summation with Histogram Encoding (SHE) [1], Thresholding with Histogram Encoding (THE) [22], Binary Local Hashing (BLH) [5] and Optimal Local Hashing (OLH) [22]. These protocols mainly focus on improving utility, often quantified in terms of the variance (i.e., Mean Squared Error – MSE), as well as optimizing computational and communication costs [23, 24] to enable efficient data collection in large-scale systems.
@@
While utility and communication costs have been the traditional focus of LDP frequency estimation protocols, recent research has shed light on their vulnerabilities in adversarial settings. For instance, re-identification risks [25, 26] have shown that adversaries can uniquely identify users within a dataset, while attribute inference attacks [27, 28, 9] have been explored in the context of iterative data collections. Furthermore, recent works have examined data reconstruction attacks [28, 26, 29], in which adversaries aim to recover the user’s original input from its obfuscated output. Among these threats, we focus specifically on data reconstruction attacks, which are particularly critical, as they represent a worst-case leakage scenario: if an attacker can recover the user’s data from a single report, other adversaries with access to side information or repeated queries can only perform better.
Figure 1: Comparison of data reconstruction attack (i.e., ASR) vs. variance (i.e., MSE) for four state-of-the-art LDP protocols (SS [20, 21], OUE [22], OLH [22], THE [22]) and our newly proposed adaptive versions (ASS, AUE, ALH, ATHE). Each subplot considers a range of privacy budgets $\varepsilon\in(2,10)$ and a fixed domain size $k=100$. Our adaptive protocols (indicated by $\circ$ markers) yield substantially lower ASR at the same or close levels of MSE when compared to their original counterparts (indicated by $\square$ markers). This improvement is reflected in the adaptive protocols’ proximity to the Pareto frontier, indicating a more favorable privacy-utility trade-off and reduced vulnerability to privacy attacks.
Contributions. In this work, we introduce a general multi-objective optimization framework for refining LDP frequency estimation protocols, enabling the joint optimization of privacy and utility under various adversarial settings. While our framework is flexible enough to incorporate multiple objectives, in this paper we focus on data reconstruction attacks [28, 26, 29] quantified by the Attacker Success Rate (ASR), and utility measured via Mean Squared Error (MSE). As aforementioned, data reconstruction attacks are particularly relevant as they directly challenge the fundamental goal of LDP: preventing an adversary from inferring a user’s value from the obfuscated output. Meanwhile, MSE serves as a widely adopted utility metric in LDP literature [30, 22, 31] due to its analytical tractability and connection to other estimation measures, such as Mean Absolute Error (MAE) and Fisher Information [19, 32] (the rationale for choosing these metrics is further discussed in Section V-B).
To demonstrate the practical benefits of our framework, we refine four state-of-the-art LDP protocols (SS, OUE, OLH and THE), by proposing adaptive versions, namely, ASS, AUE, ALH and ATHE, which achieve superior trade-off between privacy and utility compared to their traditional counterparts (e.g., see Figure 1). Additionally, we derive the expected data reconstruction attack (i.e., analytical ASR) for three additional LDP protocols (SHE, THE and a generic Unary Encoding protocol), extending prior analyses [28, 29]. Our refined protocols offer enhanced robustness against privacy attacks while maintaining practical utility levels, demonstrating the effectiveness of our framework in optimizing LDP mechanisms beyond conventional single-objective approaches. The main contributions of this paper can be summarized as follows:
@@
  • We introduce a general multi-objective optimization framework that enables the joint optimization of privacy and utility in LDP frequency estimation. Our framework is designed to assist practitioners in deploying LDP mechanisms with customized privacy-utility trade-offs, acting as a tuning layer adaptable to various systems (e.g., telemetry collection, mobile analytics).
  • We extend four state-of-the-art LDP frequency estimation protocols—SS, OUE, OLH and THE—by proposing refined versions named ASS, AUE, ALH and ATHE. These refined protocols offer a significantly better trade-off between robustness to privacy attacks and utility compared to their traditional counterparts (e.g., see Figure 1).
@@
  • We derive the analytical closed-form equation of the expected data reconstruction attack for three LDP protocols going beyond previous works [28]. These derivations are critical for evaluating and optimizing the guarantees of LDP protocols under adversarial inference scenarios.
  • We validate our proposed adaptive protocols through extensive experiments, demonstrating their effectiveness in achieving a more favorable balance between privacy (ASR) and utility (MSE) compared to existing protocols, under the same $\varepsilon$-LDP guarantees. Our results indicate that the refined protocols can substantially reduce adversarial success rates while maintaining competitive estimation accuracy.
Outline. The remainder of this paper is structured as follows. First, Section II reviews the most relevant related work, providing context for our contributions. Next, Section III defines the problem, introduces the LDP privacy model and describes the adversarial model used in this study. Section IV provides an overview of the eight LDP frequency estimation protocols analyzed in this work, along with their attack models. Section V then presents our multi-objective framework and the proposed adaptive LDP protocols that refine the original methods to enhance privacy and utility. Afterward, Section VI details the experiments conducted and presents a comprehensive analysis of the results. Finally, Section VII concludes the paper and discusses future research.
@@ II Related Work
Frequency estimation is one of the primary objective of LDP, serving as a building block for a wide range of advanced applications, such as heavy hitter estimation [5, 6, 7], frequency monitoring [3, 4, 8, 9], multidimensional queries [10, 11, 12, 13], frequent item-set mining [14, 15], and spatial density estimation [16, 17]. Numerous LDP frequency estimation protocols have been proposed in the literature [19, 20, 22, 21, 23, 31, 24], primarily focusing on minimizing estimation error, computational cost, or communication overhead. However, recent studies started to examine LDP frequency estimation mechanisms from an adversarial perspective such as data reconstruction attacks [28, 26, 29], re-identification risks [25, 26], attribute inference attacks [27, 9, 33] and poisoning attacks [34]. Our focus is on data reconstruction attacks, which allow adversaries to predict the user’s input based on the observed obfuscated output. However, unlike these prior studies that primarily highlight vulnerabilities, we propose a systematic methodology to refine existing protocols, thus reducing their susceptibility to known and emerging privacy threats.
Specifically, our work differs from the existing data reconstruction attack literature [28, 26, 29] in the following aspects. First, we formally analyze three LDP protocols’ expected ASR beyond the ones in [28]. Second, we extensively analyzed the privacy, utility, and robustness against privacy attacks of eight state-of-the-art LDP protocols. Third, we formulate a multi-objective optimization problem for LDP frequency estimation protocols instead of the single-objective one (i.e., utility-driven). This allowed us to propose four new adaptive and refined LDP protocols, which provide better trade-offs in terms of privacy, utility, and robustness against adversarial attacks.
@@ III Background
Notation. We use italic uppercase letters (e.g., $U$) to denote sets, and write $[n]=\{1,\ldots,n\}$ to represent a set of $n$ positive integers. Vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), where $\mathbf{x}_{i}$ represents the value of the $i$-th coordinate of $\mathbf{x}$. Finally, randomized mechanisms are denoted by ${\mathcal{M}}$, the input domain is denoted by ${\mathcal{X}}$, and the output domain by ${\mathcal{Y}}$. Both ${\mathcal{X}}$ and ${\mathcal{Y}}$ are discrete, in which $|{\mathcal{X}}|=k$ and $|{\mathcal{Y}}|$ depends on the randomized mechanism ${\mathcal{M}}$.
@@ III-A Problem Statement
We consider an untrusted server collecting data from a distributed group of users while preserving their privacy. Formally, there are $n$ users, with each user holding a value $x$ from a discrete domain ${\mathcal{X}}=\{1,2,\ldots,k\}$. The task is frequency estimation, in which the server aims to learn the frequencies of each value across all users, denoted as $\mathbf{f}=\{f_{i}\}_{i\in[k]}$.
@@
  • Adversary’s goal. The adversary aims to accurately infer each user’s true value $x\in{\mathcal{X}}$ based on the obfuscated output $y\in{\mathcal{Y}}$ sent to the server (i.e., data reconstruction attack).
This threat model captures a worst-case adversary (e.g., a corrupted or server itself a man-in-the-middle adversary) who observes the user’s obfuscated report and aims to infer the original input without auxiliary information, reflecting a realistic scenario in distributed or untrusted settings.
@@ III-C Pure LDP Framework
As shown in [22], the estimator in Equation (2) is unbiased (i.e., $\mathbb{E}[\hat{f}_{i}]=f_{i}$) and the variance of the estimation $\hat{f}_{i}$ is:
@@
With a sufficiently large domain size and no dominant frequency $f_{i}$, the second term of Equation (3) can be ignored. Thus, as commonly used in the LDP literature [22, 8, 11], we will consider the approximate variance, which is given by:
@@ III-D Data Reconstruction Attack on LDP
The fundamental premise of $\varepsilon$-LDP, as stated in Equation (1), is that the input to ${\mathcal{M}}$ cannot be confidently determined from its output, with the level of confidence determined by $e^{\varepsilon}$. Therefore, the user’s privacy is considered compromised if an adversary ${\mathcal{A}}$ can successfully infer the user’s original input $x\in{\mathcal{X}}$ from the obfuscated output $y\in{\mathcal{Y}}$. While various adversarial models have been proposed to exploit vulnerabilities in LDP mechanisms, including re-identification [25, 26] and poisoning [34] attacks, in this paper, and without loss of generality, we focus on data reconstruction attacks [28, 26, 29]. These attacks provide a fundamental measure of privacy leakage in LDP, as they directly quantify how well an adversary can infer the true input from the obfuscated response. Formally, the adversary’s prediction $\hat{x}$ can be defined as:
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[x\mid y]\mathrm{.}$
By applying Bayes’ theorem, we rewrite the expression as:
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\frac{\Pr[y\mid x]\cdot\Pr[x]}{\Pr[y]}\mathrm{.}$
Assuming a uniform prior distribution over $x$ (i.e., $\Pr[x]=\frac{1}{k}$, where $k=|{\mathcal{X}}|$), and noting that $\Pr[y]$ is constant for a given observation, the expression simplifies to:
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[y\mid x]\mathrm{.}$
To assess the accuracy of the attack, we will use the Attacker Success Rate (ASR) metric, which represents the probability that the adversary’s prediction $\hat{x}$ matches the input value $x$ (i.e., correctly reconstructs the users’ data):
$\text{ASR}=\Pr[\hat{x}=x]=\frac{1}{n}\cdot\sum_{j=1}^{n}\mathbb{1}\left({\hat{x}^{j}=x^{j}}\right)\mathrm{.}$
Mathematically, it is also possible to derive the expected data reconstruction attack for each protocol through formal expected value analysis [28]:
$\mathbb{E}[\text{ASR}]=\mathbb{E}[\Pr[\hat{x}=x]]\mathrm{.}$
With ASR, we can assess LDP protocols’ effectiveness in protecting privacy, uncovering vulnerabilities not evident from $\varepsilon$ alone.
IV LDP Frequency Estimation Protocols

### v3

@@ Abstract
Local Differential Privacy (LDP) offers strong privacy protection, especially in settings in which the server collecting the data is untrusted. However, designing LDP mechanisms that achieve an optimal trade-off between privacy, utility and robustness to adversarial inference and integrity attacks remains challenging. In this work, we introduce a general multi-objective optimization framework for refining LDP protocols, enabling the joint optimization of privacy and utility under various adversarial settings. While our framework is flexible to accommodate multiple privacy and security attacks as well as utility metrics, in this paper, we specifically optimize for Attacker Success Rate (ASR) under data reconstruction attack as a concrete measure of privacy leakage and Mean Squared Error (MSE) as a measure of utility. Complementarily, we evaluate integrity-oriented threats through data poisoning attacks, providing an additional adversarial perspective. More precisely, we systematically revisit these trade-offs by analyzing eight state-of-the-art LDP frequency estimation protocols and proposing refined counterparts that leverage tailored optimization techniques. Experimental results demonstrate that our proposed adaptive mechanisms consistently outperform their non-adaptive counterparts, achieving substantial reductions in ASR while preserving utility, and pushing closer to the ASR-MSE Pareto frontier. By bridging the gap between theoretical guarantees and real-world vulnerabilities, our framework enables modular and context-aware deployment of LDP mechanisms with tunable privacy-utility-attackability trade-offs.
@@ I Introduction
The local DP model has gained widespread adoption, with major technology companies integrating it into their systems to enhance user privacy. Notable examples include Google Chrome [3], Apple iOS [4] and Windows 10 [5], in which LDP protocols have been used to collect statistics while safeguarding individual data. A fundamental task under LDP guarantees is frequency estimation, which forms the basis of many advanced data analysis tasks like heavy hitter estimation [6, 7, 8], frequency monitoring [3, 5, 9, 10], multidimensional queries [11, 12, 13, 14], frequent item-set mining [15, 16] and spatial density estimation [17, 18].
@@
Due to its importance, numerous LDP frequency estimation protocols have been proposed [19], namely, Generalized Randomized Response (GRR) [20], Subset Selection (SS) [21, 22], Symmetric Unary Encoding (SUE) [3], Optimized Unary Encoding (OUE) [23], Summation with Histogram Encoding (SHE) [1], Thresholding with Histogram Encoding (THE) [23], Binary Local Hashing (BLH) [6] and Optimal Local Hashing (OLH) [23]. These protocols mainly focus on improving utility, often quantified in terms of the variance (i.e., Mean Squared Error – MSE), as well as optimizing computational and communication costs [24, 25, 26] to enable efficient data collection in large-scale systems.
@@
While utility and communication costs have been the traditional focus of LDP frequency estimation protocols, recent research has shed light on their vulnerabilities in adversarial settings. For instance, re-identification risks [27, 28] have shown that adversaries can uniquely identify users within a dataset, while attribute inference attacks [29, 30, 10] have been explored in the context of iterative data collections. Furthermore, recent works have examined data reconstruction attacks [30, 28, 31], in which adversaries aim to recover the user’s original input from its obfuscated output. Among these threats, we focus on particularly critical reconstruction attacks as they represent a worst-case leakage scenario: if an attacker can recover the user’s data from a single report, other adversaries with access to side information or repeated queries can only perform better (Examples 1 and 2).
Contributions. This paper introduces a novel attack-aware evaluation framework for analyzing and improving LDP frequency estimation protocols, designed to capture the joint trade-off between privacy leakage and utility under adversarial settings. Our approach evaluates LDP mechanisms along a unified privacy–utility–attackability trade-off curve, using the Attacker Success Rate (ASR) to quantify privacy leakage under data reconstruction attacks [30, 28, 31], and the Mean Squared Error (MSE) to measure utility [32, 23, 33]. The rationale for choosing these metrics is further discussed in Section V-B. Beyond reconstruction attacks, we further instantiate our framework to capture complementary adversarial objectives such as integrity threats through data poisoning [34], alongside ASR and MSE (see Section VI-E).
To operationalize this framework, we conduct an extensive analytical and empirical evaluation that provides new insights into the strengths and weaknesses of state-of-the-art protocols, and further enables a principled refinement of their design. In particular, we show how to tune protocol parameters through multi-objective optimization, yielding adaptive variants that better navigate the privacy-utility-attackability trade-off while preserving $\varepsilon$-LDP guarantees. Additionally, we derive the expected data reconstruction attack (i.e., analytical ASR) for three additional LDP protocols (SHE, THE and a generic Unary Encoding protocol), extending prior analyses [30, 31]. The main contributions of this paper are as follows:
@@
  • We propose a novel evaluation framework that systematically analyzes LDP protocols under the joint lens of adversarial success (targeting privacy leakage or integrity) and utility loss. Our benchmark is designed to assist practitioners in deploying LDP protocols with customized privacy-utility-attackability trade-offs, acting as a tuning layer adaptable to various systems (e.g., telemetry [5]).
@@
  • We derive new analytical closed-form equation of the expected data reconstruction attack for three LDP protocols, going beyond previous works [30]. These derivations enable principled reasoning about adversarial leakage and support optimization at the protocol-design level.
  • We conduct an extensive experimental evaluation of eight state-of-the-art LDP frequency estimation protocols, quantifying their behavior across a wide range of domain sizes and privacy regimes. Our results expose previously underexplored vulnerabilities and trade-offs.
  • Leveraging the insights from our analytical and empirical analyses, we propose refined adaptive variants (ASS, AUE, ALH, ATHE) of four well-known protocols (SS, OUE, OLH, THE). Our adaptive versions consistently achieve more favorable privacy-utility-attackability trade-offs, substantially reducing adversarial success rates while maintaining competitive estimation accuracy.
The code used in our experiments is available in our GitHub repository: https://github.com/hharcolezi/LDP_protocols_refined.
@@ II Related Work
Frequency estimation is one of the primary objective of LDP that supports many downstream applications such as heavy hitter estimation [6, 7, 8], frequency monitoring [3, 5, 9, 10], multidimensional queries [11, 12, 13, 14], frequent item-set mining [15, 16], and spatial density estimation [17, 18]. A large body of work has proposed LDP mechanisms to improve estimation accuracy, reduce communication cost, or enable scalable aggregation [20, 21, 23, 22, 24, 33, 25, 26]. While these studies have established strong theoretical and practical baselines, they primarily evaluate protocols through single-objective metrics such as MSE or communication overhead.
More recently, several works have begun to examine the adversarial dimension of LDP, including data reconstruction attacks [30, 28, 31], re-identification risks [27, 28], attribute inference [29, 10, 35], and poisoning attacks [34]. These studies revealed significant vulnerabilities in existing protocols but largely stopped at qualitative observations or empirical demonstrations, without providing a unified benchmark to compare privacy leakage and utility across mechanisms.
Our work differs by taking an experimental and analytical benchmarking perspective. We quantify vulnerabilities through privacy and integrity threats and analyze their interplay with MSE across eight representative protocols. This joint evaluation highlights protocol-dependent trade-offs for privacy-utility-attackability and offers a reproducible methodology for attack-aware benchmarking. Unlike prior analyses [30], which derive ASR for a subset of mechanisms, our study extends the analytical treatment to additional protocols and demonstrates how these insights can guide the parameter tuning of existing mechanisms, yielding refined adaptive variants.
@@ III Background
Notation. We use italic uppercase letters (e.g., $U$) to denote sets, and write $n]=\{1,\ldots,n\}$ to represent a set of $n$ positive integers. Vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), where $\mathbf{x}_{i}$ represents the value of the $i$-th coordinate of $\mathbf{x}$. Finally, randomized mechanisms are denoted by $\mathcal{M}$}, the input domain is denoted by $\mathcal{X}$}, and the output domain by $\mathcal{Y}$}. Both $\mathcal{X}$} and $\mathcal{Y}$} are discrete, in which $|{\mathcal{X}}|=k$ and $|{\mathcal{Y}}|$ depends on the randomized mechanism $\mathcal{M}$}.
@@ III-A Problem Statement
We consider an untrusted server collecting data from a distributed group of users while preserving their privacy. Formally, there are $n$ users, with each user holding a value $x$ from a discrete domain ${\mathcal{X}}=\{1,\ldots,k\}$. The task is frequency estimation, in which the server aims to learn the frequencies of each value across all users, denoted as $\mathbf{f}=\{f_{i}\}_{i\in[k]}$.
@@
  • Adversary’s goal. The adversary seeks to exploit the obfuscated reports to break the system guarantees, either by inferring users’ true values (i.e., data reconstruction) or by manipulating the aggregation process to bias the estimated frequencies (i.e., data poisoning).
This threat model captures a worst-case adversary (e.g., a curious or compromised server or a man-in-the-middle) who observes or injects obfuscated reports in an untrusted setting.
@@ III-C Pure LDP Framework
As shown in [23], the estimator in Eq. (2) is unbiased (i.e., $\mathbb{E}[\hat{f}_{i}]=f_{i}$) and the variance of the estimation $\hat{f}_{i}$ is:
@@
With a sufficiently large domain size and no dominant frequency $f_{i}$, the second term of Eq. (3) can be ignored. Thus, as commonly used in the LDP literature [23, 9, 12], we will consider the approximate variance, which is given by:
@@ III-D Data Reconstruction Attack on LDP
The fundamental premise of $\varepsilon$-LDP is that the input to $\mathcal{M}$} cannot be confidently inferred from its output, with the level of confidence bounded by $e^{\varepsilon}$. A user’s privacy is thus compromised if an adversary $\mathcal{A}$} can accurately infer the original input $x\in{\mathcal{X}}$ from the obfuscated output $y\in{\mathcal{Y}}$. To assess this residual inference risk, we focus on data reconstruction attacks [30, 28, 31], which directly quantify how well an adversary can recover the true input from the perturbed report.
Formally, the adversary infers:
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[x\mid y].$ (5)
Applying Bayes’ theorem gives
$\hat{x}=\arg\max_{x\in{\mathcal{X}}}\Pr[y\mid x]\cdot\Pr[x]\mathrm{.}$ (6)
Without loss of generality, we adopt a uniform prior, i.e., $\Pr[x]=1/k$ with $k=|\mathcal{X}|$, to obtain a protocol-intrinsic and distribution-agnostic measure of attackability. Thus, Eq. (6) reduces to
$\hat{x}=\arg\max_{x\in\mathcal{X}}\Pr[y\mid x],$ (7)
so that the adversary’s decision depends solely on the mechanism’s likelihood function. As shown in prior work [30], incorporating informative priors naturally increases the attack success rate by injecting side information, motivating our conservative distribution-independent choice.
We quantify reconstruction risk using the Attacker Success Rate (ASR), defined as the probability that an adversary correctly reconstructs a user’s original input [30]:
$\text{ASR}=\Pr[\hat{x}=x]=\frac{1}{n}\sum_{j=1}^{n}\mathbb{1}(\hat{x}^{j}=x^{j})\mathrm{,}$ (8)
which offers a more interpretable view of privacy protection, revealing vulnerabilities not captured by $\varepsilon$.
Connection to other inference attacks. Data reconstruction constitutes a fundamental inference task under LDP. Other attacks, such as attribute inference [29, 10, 35] and re-identification [27, 28], can be viewed as special cases that build upon reconstruction by leveraging correlations across attributes or external knowledge. This is similar to prior formulations [36, 37] in the machine learning community relating data reconstruction to attribute and membership inference.
Example 1 (Reconstruction as a proxy for re-identification).
Consider a user described by $d$ quasi-identifiers $x=(x^{(1)},\dots,x^{(d)})$, each independently perturbed under an $\varepsilon$-LDP mechanism. Let $\hat{x}^{(i)}$ denote the adversary’s reconstruction of attribute $x^{(i)}$ from the corresponding privatized report. The joint reconstruction success is defined as [28]:
$\mathrm{ASR}_{\mathrm{joint}}=\Pr[\hat{x}^{(1)}=x^{(1)}\wedge\ldots\wedge\hat{x}^{(d)}=x^{(d)}].$
Under the conditional independence of the reconstructions, this reduces to $\mathrm{ASR}_{\mathrm{joint}}=\prod_{i=1}^{d}\mathrm{ASR}_{i}$. Correctly reconstructing all quasi-identifiers might uniquely characterize a record within a population [38], making $\mathrm{ASR}_{\mathrm{joint}}$ a natural proxy for linkage attacks and re-identification.
Example 2 (Reconstruction as a proxy for attribute inference).
Consider an iterative data collection setting in which a user reports their value $x_{t}\in\mathcal{X}$ at each time step $t$ using an $\varepsilon$-LDP mechanism. Following standard deployments based on memoization (e.g., Microsoft’s ddBitFlipPM [5]), the same privatized output is reused across time as long as the underlying value remains unchanged. Let $\hat{x}_{t}$ denote the adversary’s reconstruction at time $t$. If reconstruction succeeds with non-negligible probability, the adversary can infer data changes by detecting transitions in the reconstructed outputs [10]:
$\Pr[\hat{x}_{t}\neq\hat{x}_{t-1}]\approx\Pr[x_{t}\neq x_{t-1}].$
As a result, successful reconstruction enables attribute inference over time, even when individual reports satisfy $\varepsilon$-LDP [35, 29]. This allows an adversary to infer sensitive behavioral changes (e.g., starting a diet or going on holiday) without necessarily recovering the exact value itself.
IV Overview of Pure LDP Frequency
Estimation Protocols